OpenStudy (anonymous):

Finish writing the equation of the line y=1/4x + ____, given the points (-12, 0) and (4, 4). Type only the part that is missing in the equation.

OpenStudy (saifoo.khan):

3

OpenStudy (anonymous):

really? ow did u get tht? show me please?

OpenStudy (saifoo.khan):

Sure.

OpenStudy (saifoo.khan):

they have already given you the slope, now just insert any set of values. y = 1x/4 + b let's take the second set. (4, 4) (4) = (4)1/4 + b 4 = 1 + b b = 4 -1 b = 3

OpenStudy (anonymous):

oh wow tht was east kool thanx hun

OpenStudy (saifoo.khan):

lol, you welcome! :D